what is the diagonal measurement of a rectangle that is 20 ft by 76 ft
Answer by jim_thompson5910(35256)

d = sqrt(L^2+W^2)

d = sqrt((76)^2+(20)^2)

d = sqrt(5776+400)

d = sqrt(6176)

d = 4*sqrt(386)

d = 78.5875308 (this is approximate)


So the diagonal is about 78.5875308 ft long.